### v3

# Rewriting Pre-Training Data Boosts LLM Performance in Math and Code

Abstract
The performance of large language models (LLMs) in program synthesis and mathematical reasoning is fundamentally limited by the quality of their pre-training corpora. We introduce two openly licensed datasets, released under the Llama 3.3 Community License, that significantly enhance LLM performance by systematically rewriting public data. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into self-contained, algorithmically efficient examples. Unlike prior methods that rely on exclusionary filtering or limited transformations, our transform-and-retain approach upgrades low-quality code, maximizing data utility. SwallowMath ($\approx$2.3 billion tokens) enhances Finemath-4+ by removing boilerplate, restoring context, and reformatting solutions into concise, step-by-step explanations. Within a fixed 50 billion token training budget, continual pre-training of Llama-3.1-8B with SwallowCode boosts pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to Stack-Edu, surpassing the baseline model’s code generation capabilities. Similarly, substituting SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. Ablation studies confirm that each pipeline stage contributes incrementally, with rewriting delivering the largest gains. All datasets, prompts, and checkpoints are publicly available, enabling reproducible research and advancing LLM pre-training for specialized domains.
1 Introduction
Figure 1: Comparison of Python-only datasets in 50 billion tokens continual pre-training setting. SwallowCode achieves the highest pass@1 on HumanEval(left) and HumanEval+(right) compared to CodeParrot-Clean, The-Stack-v1, The-Stack-v2-Smol, and Stack-Edu.
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities across diverse tasks, yet their proficiency in mathematical reasoning and program synthesis remains constrained by the quality of pre-training corpora. Existing public datasets for specialized domains, such as The-Stack-v1 and v2 for code [1, 2] and Finemath-4+ for mathematics [3], rely primarily on rule-based extraction from web crawls (e.g., CommonCrawl) [4] or model-based scoring to filter low-quality samples. However, these approaches often retain noisy, redundant, or stylistically inconsistent data, limiting their effectiveness–particularly in the growing trend of multi-stage pretraining (or mid-training) aimed at enhancing mathematical reasoning and program synthesis (e.g., OLMo 2 [5], Nemotron-H [6], and Phi-4 [7]). For instance, as shown in Figure 1, continual pre-training of Llama-3.1-8B [8] on The-Stack-v1/v2 typically maintains baseline performance on benchmarks like HumanEval and HumanEval+ but struggles to achieve significant gains due to unaddressed data quality issues. Unlike prior methods that focus on filtering or preserving original samples, we propose rewriting pre-training corpora to eliminate noise and redundancy, delivering high-quality, self-contained data that enables efficient model learning.
This paper introduces two openly licensed datasets under the Llama 3.3 Community License, designed to advance LLM capabilities in code generation and mathematical reasoning. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: sequential syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into algorithmically efficient, self-contained examples. By rewriting, rather than merely filtering, we eliminate noise and redundancy that persist in datasets like Stack-Edu, providing high-quality data that drives rapid accuracy improvements (Figure 1). SwallowMath ($\approx$2.3 billion tokens) transforms Finemath-4+ by removing boilerplate, restoring missing context, and reformatting solutions into concise and step-by-step explanations. Although our experiments focus on Python for ease of automated evaluation, the pipeline is language-agnostic: any programming language with a parseable syntax and a linter can benefit from the same treatment, broadening its applicability to diverse coding domains.
Continual pre-training of Llama-3.1-8B [8] for 50 billion tokens on mixed dataset with SwallowCode and multilingual corpora improves pass@1 by +17.0 on HumanEval and +16.1 on HumanEval+ compared to an equivalent budget using Stack-Edu [3]. Similarly, substituting Finemath-4+ with SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. To ensure the robustness of these gains, we conducted rigorous checks for test-set leakage, streaming the entire 16.1B-token SwallowCode corpus and finding no exact matches or high-similarity documents (Jaccard similarity $\geq$ 0.8) with HumanEval or HumanEval+ prompts. All datasets, prompts, and checkpoints are publicly released, enabling reproducible research. SwallowCode and SwallowMath provide a scalable framework for enhancing LLM pre-training, driving advancements in automated reasoning and software development.
2 Related work
2.1 Open code corpora
The Stack v1 [1] aggregates approximately 3TB of permissively licensed source code from public GitHub repositories. Although deduplication and license filtering are applied, boilerplate and autogenerated files are retained without language-specific refinement. The Stack v2 [2] extends this pipeline by sourcing data from the Software Heritage archive, relaxing licensing constraints, and introducing language-specific filters to enhance data quality. With a scale of approximately 900 billion tokens, it represents a significant increase in volume over v1. However, like its predecessor, v2 adopts a filtering-only approach, preserving files without semantic enhancements, which leaves inconsistent coding styles and fragmented scripts unaddressed.
2.2 Classifier-based filtering for code corpora
Recent work has shown that classifier-based filtering strategies, such as the FineWeb-Edu approach [9], can be effective for curating high-quality datasets [10]. These methods aim to improve model performance on code-related tasks by selecting semantically rich and well-documented samples from large-scale corpora. In this context, Stack-Edu [3] represents a significant effort to create a filtered variant of StarCoder2Data [2] prioritizing high-quality code. Stack-Edu begins by selecting the 15 most prevalent programming languages from StarCoder2Data, forming a subset of approximately 450 billion tokens. To assess code quality, Stack-Edu leverages Llama-3-70B-Instruct [8] to generate synthetic annotations for 500,000 code fragments, rating each on a 0–5 scale based on educational and structural quality. These annotations train language-specific classifiers based on the StarEncoder model [11], achieving F1 scores above 0.7 for most languages when applying a threshold of 3 for binary classification. The resulting dataset comprises 125 billion tokens and demonstrates improved model convergence and higher pass@1 scores on HumanEval compared to unfiltered corpora [3]. However, Stack-Edu’s exclusionary filtering strategy discards low-scoring snippets rather than rewriting or augmenting them, leaving residual issues—such as missing context or inconsistent naming conventions in the retained data.
2.3 LLM-driven pre-training corpus rewriting
Efforts to enhance code datasets using LLMs have gained traction. Austin et al. [12] explored LLM capabilities for program synthesis, finding that human feedback significantly improves code accuracy, but their work focused on synthesis rather than dataset rewriting. Cosmopedia [13] demonstrated the potential of synthetic data generation, using Mixtral-8×7B-Instruct [14] to create high-quality text corpora, though it did not address code-specific challenges.
Jain et al. [15] applied LLM-driven code transformations to instruction-tuning exemplars, focusing on variable renaming, modularization, and comment addition. While effective for fine-tuning, their approach is limited in scope and context compared to our SwallowCode pipeline. Their transformations address only a subset of stylistic improvements, whereas our Style-Guided Code Rewriting (SGCR) pipeline comprehensively enforces ten criteria from the Google Python Style Guide, including descriptive variable naming, effective type annotations, modular function design, error handling, and readability-focused formatting (see Section 3.3.1). Furthermore, our Self-Contained Optimization Rewriting (SCOR) pipeline introduces semantic enhancements—ensuring self-containment by resolving dependencies, optimizing algorithms, and transforming trivial snippets into instructive examples—which are absent in Jain et al.’s work (see Section 3.3.2). Additionally, Jain et al. target instruction-tuning data, a smaller and more curated dataset, whereas SwallowCode systematically rewrites large-scale pre-training corpora, a more challenging task due to the diversity and volume of code samples. Our pipeline integrates rigorous preprocessing with syntax error and linter-based filtering to ensure high-quality inputs for rewriting.
SwallowCode’s transform-and-retain paradigm upgrades low-quality code rather than discarding it, addressing limitations of existing datasets like The Stack v1 [1] and v2 [2], and Stack-Edu [3], which rely on filtering. By combining filtering, SGCR, and SCOR, SwallowCode produces a high-quality corpus that significantly improves performance on HumanEval and HumanEval+ benchmarks, advancing the state-of-the-art in code dataset curation.
2.4 Synthetic Data Generation for Code
Recent approaches, such as Magpie [16], leverage high-performing LLMs to generate synthetic instruction-tuning datasets from prompts and desired characteristics without relying on existing data. However, we opted against generating synthetic code data from scratch for SwallowCode due to two key limitations. First, previous work [17] demonstrates that low-diversity synthetic data restricts LLM performance. Second, achieving high diversity in synthetic code datasets requires diverse topics and keywords as seeds, as seen in Nemotron-4 340B’s synthetic instruction data [18]. For code, defining such seeds (e.g., varied algorithmic paradigms or problem domains) remains an open challenge, as no established methodology ensures sufficient diversity across programming tasks. Instead, our approach leverages high-quality code from The-Stack-v2, filtered for syntactic and stylistic rigor (Section 3.2), and applies LLM-driven rewriting to enhance quality while preserving the inherent diversity of real-world code. This transform-and-retain strategy maximizes data utility and avoids the risks of synthetic data homogeneity.
3 Construction of the code corpus
Figure 2: Four-stage pipeline for constructing SwallowCode: (1) syntax filtering to remove invalid Python code, (2) linter-based filtering using pylint to enforce coding standards, and (3–4) two-stage LLM rewriting with Style-Guided Code Rewriting (SGCR), which enforces consistent style and readability, and Self-Contained Optimization Rewriting (SCOR), which ensures self-containment and optimizes algorithms for efficiency.
The development of SwallowCode is driven by an empirical and exploratory approach, informed by data ablation experiments evaluating each stage of the data processing pipeline, as illustrated in Figure 2. Specifically, we continually pre-trained the baseline model with the dataset only differing code text subset: before and after applying specific stage in the pipeline, and decide if we should adopt or discard the stage based on the evaluation results. In this section, we present the experimental results and describe the design choices shaping the pipeline. Please refer to Appendix H.1 for detailed results. To ensure full reproducibility, all code, configurations, and supporting materials are publicly available at https://github.com/rioyokotalab/swallow-code-math.
3.1 Experimental setup
To evaluate the impact of each design choice in our data processing pipeline, we conduct systematic data ablation studies. Each ablation trains a model that differs only in the target pre-training dataset, holding all other factors constant, including model architecture, parameter count, non-target data, total token budget, and hyperparameters. Specifically, we perform continual pre-training starting from Llama-3.1-8B, using a total of approximately 50 billion tokens, a maximum sequence length of 8,192, a global batch size of approximately 4 million tokens, and the Llama-3 tokenizer. The target dataset is processed with less than one epoch within each ablation. We evaluate model checkpoints approximately every 10 billion tokens using ten downstream benchmarks.
Continual pre-training is conducted using Megatron-LM (version core_r0.9.0) [19]. For evaluation, we utilize the BigCodeBench [20] and lm-evaluation-harness [21] on the following benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWINO [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31]. The effectiveness of the code corpus is specifically assessed using HumanEval and HumanEval+. The pre-training data mixture consists of 84% multilingual text and 16% code. Detailed proportions and data source are provided in Appendix A.4.1, with comprehensive training hyperparameters reported in Appendix A.1. All ablation models, associated checkpoints, and supporting materials are publicly available in our Hugging Face repository to ensure full reproducibility.
3.2 Filtering
To construct the SwallowCode corpus, a high-quality Python code dataset, we implement a rigorous filtering pipeline starting with the-stack-v2-train-smol-ids [2] as the base dataset. The filtering process is critical to ensure that only syntactically correct and well-structured code is retained, enhancing downstream performance in code generation tasks. We focus exclusively on Python code to maintain consistency across ablation studies and enable fair comparisons with existing public corpora. Our pipeline employs two key filtering techniques, syntax error filtering and linter-based filtering, that significantly improve code quality. We also evaluated LLM-based scoring in ablation experiments but did not adopt it in the final pipeline due to its limited performance gains relative to computational cost. Figure 3 summarizes the performance of these methods on the HumanEval and HumanEval+ benchmarks, demonstrating the impact of our filtering strategy.
Figure 3: Performance comparison of filtering methods on HumanEval (left) and HumanEval+ (right) benchmarks. The original dataset (Stack v2) is compared against datasets processed with syntax error filtering, linter-based filtering, and LLM-based scoring (evaluated in ablation studies but not adopted). Syntax and linter-based filtering enhance code generation performance, while LLM-based scoring provides marginal gains, considering the computational cost.
3.2.1 Syntax error filtering
Despite the heuristic filtering in the BigCode project, the-stack-v2-train-smol-ids [2] includes Python code samples with invalid syntax according to Python 3.10 specifications. To address this, we apply syntax error filtering by compiling each code sample using Python’s built-in compile() function, discarding any samples that fail to compile. This process reduces the dataset from approximately 41 million to 37 million samples, a 9.7% reduction. As shown in Figure 3, removing syntactically invalid samples improves performance on HumanEval and HumanEval+  [^1] The performance trajectory of The Stack v2 exhibits an initial decline followed by a recovery, consistent with forgetting and subsequent adaptation observed in prior continual pretraining studies [32]. This behavior is not particularly noteworthy.. Consequently, we adopt syntax-error filtering as a standard step in all subsequent experiments.
3.2.2 Linter-based filtering
Beyond syntactic correctness, code quality depends on adherence to coding standards. Many samples in the initial dataset exhibit poor structure, generating numerous warnings when analyzed by static analysis tools. We employ Pylint, a widely-used Python linter, to enforce a quality threshold, excluding samples with scores below 7.0 on a 0–10 scale based on rule violations. Additionally, we penalize overly verbose comments using a custom heuristic scoring algorithm (detailed in Appendix B). This step reduces the dataset from 36.7 million to 24.1 million samples, a 34.3% reduction. Figure 3 illustrates the performance gains of more than 1 point in HumanEval and HumanEval + achieved by linter-based filtering. Consequently, we adopt linter-based filtering, with a threshold of 7.0, as a standard step in all subsequent experiments.
3.2.3 LLM-based scoring filtering
Recent approaches leverage large language models (LLMs) to generate synthetic annotations for training quality classifiers, which are then used to filter web-scale corpora by retaining high-quality samples (e.g., FineWeb [9], Stack-Edu [3]). Instead of training a separate classifier, we directly prompt Llama-3.3-70B-Instruct [8] to evaluate each Python code snippet on a scale of 0–10, based on ten criteria, including code readability, modularity, and adherence to naming conventions, derived from the Google Python Style Guide. The detailed scoring prompt and the distribution of the quality scores are provided in Appendix C.
We exclude samples scoring below 6, retaining only those deemed sufficiently high-quality, and use this filtered subset alongside multilingual data for continual pre-training in our ablation studies. As shown in Figure 3, LLM-based filtering yields modest improvements (less than 1 point) over linter-based filtering on the HumanEval and HumanEval+ benchmarks.
Given these limited gains, we compare LLM-based scoring to our LLM-driven rewriting pipeline, which refines code snippets by enhancing clarity and correctness (Section 3.3). Comparing Figure 3 and Figure 4, although the rewriting pipeline requires 1.22 times the computational resources of LLM-based scoring (a 22% increase), it achieves significantly greater performance gains on HumanEval and HumanEval+ (detailed in Appendix F). Consequently, we do not incorporate LLM-based scoring in subsequent experiments, which favors the more effective and cost-efficient rewriting approach.
3.3 LLM-driven rewriting
Figure 4: Performance comparison of LLM-driven rewriting steps on HumanEval (left) and HumanEval+ (right) benchmarks. The pre-rewriting (syntax-error and linter-based filtering) is compared against SGCR and SCOR. SGCR improves performance by over 7-9 points, while SCOR, applied after SGCR, further enhances scores by over 5-6 points, demonstrating the effectiveness of stylistic and semantic rewrites in the SwallowCode pipeline.
Recent studies highlight the potential of large language models (LLMs) to transform training data, enhancing the performance of instruction-tuned models. Jain et al. [15] proposed three transformations—variable renaming, modularization, and plan annotation—for code cleaning, demonstrating their effectiveness for instruction tuning. However, their approach is limited to curated instruction-tuning data and a narrow set of stylistic improvements, lacking semantic optimizations or preprocessing integration.
To address these limitations, we developed the SwallowCode pipeline, which systematically rewrites large-scale pre-training corpora using two complementary LLM-driven processes: Style-Guided Code Rewriting (SGCR) and Self-Contained Optimization Rewriting (SCOR). SGCR revises Python code snippets based on the Google Python Style Guide  [^2] https://google.github.io/styleguide/pyguide.html, enforcing stylistic improvements like clear naming and modular design (Section 3.3.1). SCOR extends SGCR by ensuring self-containment and applying semantic optimizations, such as efficient algorithms and instructive examples (Section 3.3.2). To clarify, SwallowCode retains the code-only format of The-stack-v2-train-smol-ids [2], consisting solely of Python snippets without the text-code pair structure typical of instruction-tuning datasets. The rewriting pipeline, powered by Llama-3.3-70B-Instruct, enhances code quality through stylistic and semantic transformations without introducing instructional prompts or responses. Thus, performance gains on HumanEval and HumanEval+ (Section 3.4) stem from improved data quality, not from distillation of instruction tuning capabilities.
To illustrate the scope of these transformations, Table 1 compares the coverage of SGCR and SCOR against the approach of Jain et al. [15]. While SGCR addresses stylistic criteria like type hints, error handling, and docstrings, SCOR introduces semantic enhancements, including self-containment and optimization (algorithm and data structure), broadening the scope of the SwallowCode pipeline.
Table 1: Comparison of code transformation coverage for Jain et al. [15], SGCR, and SCOR.
Criterion Jain et al. SGCR SCOR
Variable Renaming $\checkmark$ $\checkmark$ $\times$
Modularization $\checkmark$ $\checkmark$ $\times$
Comments $\checkmark$ $\checkmark$ $\times$
Type Hint $\times$ $\checkmark$ $\times$
Error Handling $\times$ $\checkmark$ $\times$
Docstring $\times$ $\checkmark$ $\times$
Self-Contained $\times$ $\times$ $\checkmark$
Optimized $\times$ $\times$ $\checkmark$
3.3.1 SGCR: Style-Guided Code Rewriting
SGCR enhances code readability by adding docstrings and type hints, unifying variable reassignment patterns, and standardizing function and class names in accordance with the Google Python Style Guide. Compared to the pre-rewriting (syntax-error and linter-based filtering), SGCR achieves improvements over 7-9 points on HumanEval and HumanEval+ as shown Figure 4. We also evaluate SGCR applied directly to the raw the-stack-v2-train-smol-ids corpus versus SGCR applied after syntax-error and linter-based filtering. As illustrated in Figure 4, the SGCR pipeline with syntax and linter filtering outperforms direct SGCR by 0.4-2.1 points on downstream code-generation benchmarks. Consequently, we adopt the pipeline that syntax and linter-based filtering prior to SGCR in all subsequent experiments.
Ablation studies reveal that SGCR significantly improves HumanEval and HumanEval+ scores but results in an approximate 10-point decrease on the MBPP benchmark [12] (detailed in Appendix G). Analysis indicates that MBPP’s solutions often use non-standard function and class names, and SGCR’s automated renaming introduces function name mismatches with MBPP’s unit tests, leading to “undefined” errors during evaluation. The identified mismatches obscure the model’s true code-generation capabilities, motivating us to exclude MBPP from our evaluation benchmarks across all experiments.
3.3.2 SCOR: Self-Contained Optimization Rewriting
Although SGCR ensures adherence to stylistic criteria, it does not modify the program semantics. Manual observation of models trained on SGCR-processed data reveals three recurring issues: (i) missing dependencies, where models attempt to import non-existent libraries or call undefined functions, causing runtime errors; (ii) inefficient algorithms, such as using naive recursion or quadratic-order algorithms for problems that linear-order or dynamic-programming solutions; and (iii) trivial snippets, such as code that merely prints constants or performs basic arithmetic, offering minimal training value.
To address these limitations while preserving SGCR’s stylistic improvements, we introduce Self-Contained Optimization Rewriting (SCOR). Guided by a ten-rule prompt (detailed in Appendix D.2), SCOR rewrites each snippet to ensure self-containment by inlining or satisfying external dependencies, replaces inefficient algorithms with more computationally efficient alternatives, and transforms trivial code into meaningful, executable examples. As illustrated in Figure 4, SCOR improves HumanEval and HumanEval+ scores by over 5-6 points compared to SGCR. These results underscore the importance of semantic-level rewrites beyond stylistic enhancements, establishing SCOR as the final stage of the SwallowCode construction pipeline. We did not conduct an ablation experiment evaluating SCOR in isolation without SGCR. However, prompt validation experiments with Llama-3.3-70B-Instruct indicated that simultaneously applying SGCR and SCOR often reduced the quality of the rewritten code due to the challenges LLMs face in balancing multiple objectives. This led to the adoption of a two-stage SGCR and SCOR approach, with the isolated SCOR evaluation deferred to future work.
Compared to the approximately 1-2 points performance gain achieved during the filtering stages (green line in Figure 4), the rewriting stages using SGCR and SCOR led to a total performance improvement of 14 points—7-9 points from SGCR and 5-6 points from SCOR. This highlights the significant potential for enhancing dataset curation by incorporating LLM-driven rewriting approach.
3.4 The Final SwallowCode dataset
Applying the complete pipeline, including syntax error filtering, Pylint-based filtering, Style-Guided Code Rewriting (SGCR), and Self-Contained Optimization Rewriting (SCOR), to the-stack-v2-train-smol-ids produces the SwallowCode corpus, comprising 16.1 billion tokens. All intermediate artifacts, including non-optimal variants, are publicly available to support future research efforts.
Comparison with existing corpora
Figure 1 compares SwallowCode with several widely used open code datasets: CodeParrot-Clean (12.8 billion tokens) [^3] https://huggingface.co/datasets/codeparrot/codeparrot-clean, The Stack v1 (98.2 billion tokens) [1], The Stack v2-Smol (36.1 billion tokens) [2], and Stack-Edu (17.9 billion tokens) [3]. For a fair comparison, we extract only the Python subsets of each corpus and, following the protocol outlined in Section 3.1, allocate 16% (8 billion) code tokens within a 50 billion-token mixed batch, ensuring each sample is processed no more than once. SwallowCode outperforms all comparable publicly available corpora on the HumanEval (pass@1) and HumanEval+ (pass@1) benchmarks, demonstrating the effectiveness of our pipeline design. Detailed results are provided in Appendix H.1.
Language generality
While our experiments focus on Python to facilitate automated evaluation, the pipeline’s components are language-agnostic, requiring only (i) static syntax checking and (ii) the availability of a linter or style tool. Thus, the methodology can be readily adapted to other programming languages with minimal modifications.
4 Construction of the math corpus
Section 3 demonstrated that LLM-driven rewriting significantly boosts coding performance. To evaluate the transferability of this approach, we apply a tailored rewriting pipeline to mathematical data. We select finemath-4+ [3], a high-quality, publicly available math corpus, as the starting point and process it through a rewriting pipeline. According to the evaluation in finemath [3], it outperforms other open mathematical datasets, including OpenWebMath [4], InfiMM-Math [33], and finemath-3+, in terms of the performance on benchmarks such as GSM8K and MATH. Given its reported superior performance and public availability, we adopt finemath-4+ as the foundation corpus for constructing our mathematical corpus to maximize the effectiveness of our rewriting approach.
4.1 Experimental setup
We adhere to the protocol outlined in Section 3.1, performing continual pre-training of Llama-3.1-8B for approximately 50 billion tokens, varying only the target math corpus. The evaluation benchmarks mirror those of Section 3.1, with HumanEval+ replaced by the MATH dataset [34], with GSM8K and MATH as the primary math-focused benchmarks. The pre-training mixture comprises 82.2% multilingual text, 13.0% code, and 4.79% math; detailed proportions and data sources are provided in Appendix A.4.2. The complete hyperparameters are listed in the Appendix A.1.
4.2 LLM-driven rewriting
Figure 5: Performance gains from LLM-driven rewriting of finemath-4+. The rewritten corpus improves GSM8K(left) by 12.4 points and MATH(right) by 7.6 points.
The Finemath‑4+ corpus [3] is a collection of documents in which snippets of mathematical text are embedded within passages that are otherwise unrelated to mathematics. In addition, the mathematical content ranges widely, from elementary arithmetic to advanced topics. This heterogeneity renders rule-based filtering challenging, as it struggles to distinguish relevant mathematical content from irrelevant artifacts. To address this, we design an LLM-driven rewriting pipeline using Llama-3.3-70B-Instruct, which not only cleans and refines the data but also enhances its quality for mathematical reasoning tasks.
The rewriting prompt instructs the model to: (1) remove residual web headers, footers, and privacy notices; (2) eliminate irrelevant metadata, such as question and answer timestamps; (3) restore missing context in incomplete questions or answers; (4) rewrite derivation steps to be concise yet comprehensive; and (5) provide clear, step-by-step solutions. Steps (1) and (2) are analogous to the syntax error and linter-based filtering applied to SwallowCode (Section 3), addressing inappropriate content that rule-based methods alone could not effectively filter. Steps (3) through (5) parallel the self-containment and style enhancements of the code rewriting pipeline, adapting them to the mathematical domain. The complete prompt is provided in Appendix E.
As shown in Figure 5, the rewritten corpus yields substantial improvements: 12.4 points on GSM8K and 7.6 points on MATH. These results demonstrate that LLM-driven rewriting, while tailored to the unique characteristics of mathematical data, successfully enhances the already high-quality finemath-4+ corpus. This confirms the generalizability of our rewriting approach beyond code, offering a robust method for improving open-domain datasets for mathematical reasoning.
5 Limitations
While SwallowCode and SwallowMath significantly improve code generation and mathematical reasoning performance of LLMs, several limitations should be noted. First, the rewriting pipelines may preserve biases present in the source data. For example, the-stack-v2-train-smol-ids may over-represent certain coding patterns (e.g., recursive algorithms over iterative ones), and Finemath-4+ may favor specific problem types (e.g., algebra over geometry). Additionally, as the rewriting process relies on Llama-3.3-70B-Instruct, the resulting datasets may reflect this model’s preferences, such as favoring certain variable naming conventions or solution strategies, potentially limiting generalizability. Second, our evaluations are confined to continual pre-training with a 50 billion token budget, as detailed in Section 3.1, to ensure controlled and reproducible experiments within computational constraints. The impact of extending pre-training beyond this budget remains unexplored, and performance trends at larger scales may differ, particularly for tasks requiring extensive training data. Third, although the SwallowCode pipeline is designed to be language-agnostic, requiring only static syntax checking and linter tools, our experiments focus exclusively on Python to facilitate automated evaluation. Empirical validation for other languages (for example, Java, C++) was not feasible due to resource constraints, limiting evidence of the pipeline’s broader applicability.
6 Conclusion
In this work, we present SwallowCode and SwallowMath, two pre-training datasets released under the Llama 3.3 Community License, designed to enhance large language model (LLM) capabilities in program synthesis and mathematical reasoning. SwallowCode, spanning approximately 16.1 billion tokens, refines Python code from The-Stack-v2 using a transform-and-retain pipeline that includes syntax validation, style-guided rewriting, and self-contained optimization. SwallowMath, with approximately 2.3 billion tokens, upgrades Finemath-4+ by reformatting mathematical content into concise, context-rich solutions. Unlike traditional approaches that discard low-quality data, our method systematically enhances it, maximizing utility and setting a new standard for dataset curation.
SwallowCode and SwallowMath achieve state-of-the-art performance among publicly available datasets for code and math reasoning, demonstrating the efficacy of our data refinement strategies. The language-agnostic design of our pipeline, which requires only parseable syntax and a linter, suggests its potential applicability to other programming languages. To foster reproducible research and drive advancements in LLM pretraining, we release SwallowCode, SwallowMath, all associated prompts, and model checkpoints. These resources establish a scalable framework for high-quality data curation, paving the way for enhanced automated reasoning and software development.
Appendix A Detailed setup for data ablation experiments
This section provides details on the training hyperparameters, training library versions, training environments, distributed training settings, and data mixture used in the dataset ablation experiments described in Sections 3 and 4.
A.1 Training hyperparameters
We performed continual pre-training from Llama-3.1-8B  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B using approximately 50 billion tokens. As shown in Table 2, the model architecture and tokenizer are identical to those of Llama-3.1-8B. The training hyperparameters are detailed in Table 3.
Table 2: Training model’s architecture
Hyperparameter Value
Architecture Llama-3
Hidden size 4 096
FFN hidden size 14 336
Number of layers 32
Number of attention heads 32
Number of key/value heads 8
Sequence length 8 192
Normalization RMSNorm
RMSNorm epsilon $1.0\times 10^{-5}$
RoPE base 500000
Attention dropout 0.0
Hidden dropout 0.0
Tokenizer Llama-3 tokenizer
Table 3: Training hyperparameters
Hyperparameter Value
Adam beta1 0.9
Adam beta2 0.95
Adam epsilon $1.0\times 10^{-8}$
Gradient clipping 1.0
Weight Decay 0.1
Learning rate (max) $2.5\times 10^{-5}$
Learning rate (min) $2.5\times 10^{-6}$
Warmup steps 1000
Warmup style linear
Decay style cosine
A.2 Training environment
We utilized the TSUBAME [^5] https://www.t4.cii.isct.ac.jp/docs/handbook.en/ supercomputer of the Institute of Science Tokyo, for training. We employed mixed precision (bfloat16) and used multiple NVIDIA H100 nodes for distributed parallel training. Each node is equipped with four NVIDIA H100 94GB GPUs, and the nodes are interconnected via InfiniBand NDR200.
We conducted continual pre-training with libraries shown in Table 4.
Table 4: Training library versions
Component / Library Version
Training library Megatron‑LM
mcore 0.9.0
CUDA Toolkit 12.4
cuDNN 9.1.0
NCCL 2.21.5
HPC‑X 2.17.1
ninja $1.1$
PyTorch 2.5.0
TransformerEngine 1.12
A.3 Distributed training settings
Training large language models on a single GPU is challenging due to both GPU memory constraints and the time required for training. In terms of GPU memory, even using the latest H100 80GB, it is difficult to train the 8B model used in this study. Moreover, even if the model parameters, gradients, and optimizer states could fit on a single GPU, training on a single GPU would require an unrealistic amount of time to complete. Therefore, in this study, we adopted distributed parallel training, combining data parallelism and model parallelism. We conducted all ablation experiments with the distributed setting shown in Table 5.
Table 5: Distributed training setting for ablation experiments
Hyperparameter Value
Data Parallelism (DP) 32
Tensor Parallelism (TP) 2
Context Parallelism (CP) 1
Pipeline Parallelism (PP) 1
Micro batch size 2
Global batch size 512
Sequence Parallelism true
Distributed optimizer true
Tensor Parallelism Communication Overlap true
A.4 Data mixture for ablation experiments
The SwallowProject [^6] https://swallow-llm.github.io/index.en.html, a collaborative effort between the Institute of Science Tokyo, Japan, and the National Institute of Advanced Industrial Science and Technology (AIST), aims to develop open-source large language models (LLMs) with strong capabilities in both Japanese and English. A key challenge in continual pre-training from high-performing models like Llama-3.1-8B is mitigating catastrophic forgetting, particularly in maintaining or improving mathematical reasoning and code generation performance. To address this, our ablation experiments were designed to enhance Llama-3.1-8B’s performance on HumanEval, HumanEval+, GSM8K, and MATH while incorporating a multilingual dataset predominantly composed of Japanese and English text. This approach led to the development of SwallowCode and SwallowMath, as detailed in Sections 3 and 4. The data mixture reflects a high proportion of Japanese text, consistent with our project’s focus on bilingual proficiency. However, as described in Sections 3.1 and 4.1, all ablation experiments maintain identical settings except for the target code or math dataset, ensuring a fully controlled experimental design.
This continual pre-training strategy aligns with established practices in high-quality LLM development, as seen in models like OLMo-2 [5] and Nemotron-H [6]. Specifically, adopting a staged pre-training approach, as observed in OLMo-2 and Nemotron-H, where later training phases leverage high-quality multilingual text alongside specialized math and code datasets to enhance LLM capabilities, ensures that our ablation experiments align with realistic LLM training scenarios.
A.4.1 Code ablation data mixture
Figure 6: Data ratio for SwallowCode ablation experiments.
The training dataset for the code ablation experiments comprises approximately 50 billion tokens. The distribution of the dataset components is illustrated in Figure 6, with the following components and their respective token counts. Note that the Target Code Data varies depending on the specific ablation experiment conducted.
A.4.2 Math ablation data mixture
Figure 7: Data ratio for SwallowMath ablation experiments.
The training dataset for the math ablation experiments also consists of approximately 50 billion tokens. The composition of the dataset is shown in Figure 7, with the following components and their respective token counts. Note that the Target Math Data varies depending on the specific ablation experiment conducted.
Appendix B Code linting filtering
A threshold of 7.0 balances code quality and dataset size, while the comment penalty reduces overly verbose or non-functional scripts. As described in Section 3.2.2, the code for performing linter filtering is publicly available at https://github.com/rioyokotalab/swallow-code-math. The relevant code is also provided below.
When conducting linting with pylint, warnings and errors dependent on the linting environment, such as import errors, are excluded using the --disable option. Additionally, some files with Python extensions primarily contain comments with textual content and have minimal script functionality. To address this, as discussed in Section 3.2.2, we introduced a mechanism that imposes a heuristic penalty based on the proportion of comments to filter out such files.
1def check_comment_ratio(code: str):
2 total_lines = 0
3 comment_lines = 0
4
5 try:
6 tokens = tokenize.generate_tokens(StringIO(code).readline)
7 for token_type, _, _, _, _ in tokens:
8 total_lines += 1
9 if token_type == tokenize.COMMENT:
10 comment_lines += 1
11
12 except tokenize.TokenError as e:
13 print(f"Token error encountered: {str(e)}")
14 return 0
15 except IndentationError as e:
16 print(f"indentation error encountered {str(e)}")
17 return 0
18
19 if total_lines == 0:
20 return 0
21
22 return comment_lines / total_lines
23
24
25def apply_comment_penalty(score: float, comment_ratio: float) -> float:
26 if comment_ratio == 1.0:
27 return 0.0
28 elif comment_ratio > 0:
29 penalty_factor = 1 - comment_ratio
30 score *= penalty_factor
31 return score
32
33
34def check_code_quality(code: str):
35 with tempfile.NamedTemporaryFile(delete=False, suffix=".py") as temp_file:
36 temp_file.write(code.encode())
37 temp_file.flush()
38
39 result = subprocess.run(
40 ["pylint", "--persistent=n","--disable=E0401,C0114,C0301,C0103,C0116,C0411,R0903,W0511,C0412", temp_file.name],
41 capture_output=True,
42 text=True,
43 )
44
45 pylint_output = result.stdout
46 score = None
47
[…]
49 if "Your code has been rated at" in line:
50 score = float(line.split("/")[0].split()[-1])
51
52 comment_ratio = check_comment_ratio(code)
53
54 if score is not None:
55 score = apply_comment_penalty(score, comment_ratio)
56
57 return score, pylint_output
Appendix C Code LLM-based scoring
As described in Section 3.2.3, this section presents the prompt used for LLM-based scoring. The prompt was provided to Llama-3.3-70B-Instruct to evaluate code quality. The scoring criteria were developed with reference to the Google Python Style Guide  [^13] https://google.github.io/styleguide/pyguide.html. The actual code pipeline is available at https://github.com/rioyokotalab/swallow-code-math. The dataset created through this process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
Prompt Used for Code Quality Evaluation You are a smart software engineer. Please evaluate the following code on a scale of 1 to 10 based on the following criteria:
1. Are variable names descriptive and consistent with naming conventions?
2. Are comments and docstrings appropriately written to explain the purpose and functionality of the code?
3. Are type annotations used effectively where applicable?
4. Are functions appropriately modularized, with well-defined responsibilities and clear separation of concerns?
5. Are variables’ lifetimes intentionally managed, avoiding frequent reassignment or overly long scopes?
6. Is error handling implemented appropriately where necessary?
7. Is the code properly indented and follows standard formatting guidelines?
8. Do comments provide context and rationale, rather than merely describing what the code does?
9. Are functions and classes designed with clear, single responsibilities?
10. Is the code formatted in a way that enhances readability?
Figure 8: Distribution of quality scores assigned by Llama-3.3-70B-Instruct to the Python code corpus. The dashed line at 6 denotes the filtering threshold. Notably, only even scores appear, an emergent property of how Llama-3.3-70B-Instruct interpreted the scoring prompts.
Figure 8 illustrates the distribution of the quality scores assigned.
Appendix D Code LLM rewriting
As described in Section 3.3, this section presents the prompts used for LLM-based rewriting, specifically for Style-Guided Code Rewriting (SGCR) and Self-Contained Optimization Rewriting (SCOR). Each prompt was provided to Llama-3.3-70B-Instruct to perform data rewriting.
D.1 Style-Guided Code Rewriting (SGCR)
The prompt used for SGCR is provided below. The dataset created through this rewriting process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
Prompt Used for SGCR You are a smart software engineer. Please evaluate the following code on a scale of 1 to 10 based on the following criteria:
1. Are variable names descriptive and consistent with naming conventions?
2. Are comments and docstrings appropriately written to explain the purpose and functionality of the code?
3. Are type annotations used effectively where applicable?
4. Are functions appropriately modularized, with well-defined responsibilities and clear separation of concerns?
5. Are variables’ lifetimes intentionally managed, avoiding frequent reassignment or overly long scopes?
6. Is error handling implemented appropriately where necessary?
7. Is the code properly indented and follows standard formatting guidelines?
8. Do comments provide context and rationale, rather than merely describing what the code does?
9. Are functions and classes designed with clear, single responsibilities?
10. Is the code formatted in a way that enhances readability?
And provide suggestions for improvement based on the evaluation criteria. You can also provide an improved version of the code in the following style:
### Evaluation: 7 ### Suggestions: Provide specific, actionable suggestions to improve the code based on the evaluation criteria.
### Improved Code: Provide a revised version of the code incorporating the suggested improvements.
‘‘‘python
def improved_function(arg1: int, arg2: str) -> str:
# Your improved code here
pass
D.2 Self-Contained Optimization Rewriting (SCOR).
The prompt used for SCOR is provided below. The dataset created through this rewriting process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
Prompt Used for SCOR You are a smart software engineer. Please change a given code into self-contained and well-structured code following the below best practices and pythonic way.
1. Use meaningful variable and function names.
2. Write a clear and concise docstring for the function.
3. Use type hints for the function signature.
4. Write a clear and concise comment for the code block.
5. Ensure the code is self-contained and does not depend on external variables.
6. Ensure the code is well-structured and easy to read.
7. Ensure the code is free of errors and runs correctly.
8. Ensure the code is optimized and does not have redundant operations.
9. Ensure the algorithm and data structures are efficient and concise.
If given code is not self-contained or too simple, please change it to a more educational and useful code.
Appendix E Math LLM rewriting
As described in Section 4.2, this section presents the prompt used for LLM rewriting in the construction of SwallowMath. The prompt consists of five components: (1) remove residual web headers, footers, and privacy notices; (2) delete extraneous metadata such as question and answer timestamps; (3) fill in missing context when either the question or answer is incomplete; (4) rewrite explanations to be concise yet information-dense; and (5) present a clear step-by-step solution. Steps (1)–(2) parallel our syntax-error and linter filtering for code, while steps (3)–(5) correspond to the self-containment and style rewrites used in SwallowCode.
Prompt Used for Math Rewriting You are an intelligent math tutor. You are given the following math problem and answer with some unnecessary parts. Please remove the unneeded parts of the questions. For example, the date of the question submitted, the answer date, the privacy policy, the footer, the header, etc., should be removed. However, please keep the main question and answer.
If questions or answers lack some information or are not elaborate, please make them more informative and easy to understand. If needed, please add more detail about the step-by-step calculation process.
Appendix F Computational cost
This section quantifies the computational cost, measured in GPU hours (H100), required for the LLM scoring, LLM rewriting, and ablation experiments conducted in this study. These measurements are based on empirical data. By analyzing the computational cost, we aim to elucidate the relationship between the resource demands of the data cleaning pipeline and the resulting improvements in downstream task performance for each method. Furthermore, by providing an estimate of the computational resources needed to reproduce our results, we seek to facilitate future research.
F.1 Computational cost of LLM scoring
The data synthesis process utilized vLLM 0.7.2 and PyTorch 2.5.1. The dataset was synthesized with a global batch size of 2048 and tensor parallelism of 4. Data generation was performed using four H100 (94 GB) GPUs per job. With an input processing speed of approximately 2000 tokens/s and an output generation speed of approximately 3000 tokens/s, and given an average input dataset length of 836 tokens, an average output dataset length of 1271 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.2.3 consumed a total of 19,477 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
F.2 Computational cost of LLM rewriting
The synthesis process for LLM rewriting also employed vLLM 0.7.2 and PyTorch 2.5.1, with a global batch size of 2048 and tensor parallelism of 4. Data generation was conducted using 4 H100 (94GB) GPUs per job. The input processing speed was approximately 2000 tokens/s, and the output generation speed was approximately 3000 tokens/s. Given an average input dataset length of 836 tokens, an average output dataset length of 1819 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.3.1 consumed a total of 23,703 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
F.3 Computational cost of continual pre-training data ablation experiments
The ablation experiments described in Sections 3.1 and 4.1 were conducted using 64 H100 (94GB) GPUs for 24.7 hours per experiment. Thus, each experiment consumed 1,580 H100 GPU hours. With a total of 15 experiments (13 code ablation experiments and two math ablation experiments), the overall computational cost amounted to 23,700 H100 GPU hours. These experiments achieved a training throughput of 530 TFLOP/s/GPU and 590,000 tokens/s, as derived from the FLOP/s formula presented in the Megatron-LM paper [37]. This corresponds to approximately 53.5% of the peak BF16 Tensor Core performance of the H100.
Appendix G MBPP
As discussed in Section 3.3.1, the MBPP dataset  [^14] https://github.com/google-research/google-research/blob/master/mbpp/mbpp.jsonl contains Python functions with naming conventions that deviate from standard Python style guidelines. For example, the following code snippet uses camelCase instead of the recommended snake_case:
1def is_Power_Of_Two(x):
2 return x and (not(x & (x - 1)))
3def differ_At_One_Bit_Pos(a, b):
4 return is_Power_Of_Two(a ^ b)
As described in Section 3.3.1, Style-Guided Code Rewriting (SGCR) rewrites code to conform to Python’s naming conventions  [^15] https://peps.python.org/pep-0008/#descriptive-naming-styles, specifically enforcing snake_case for function names. Consequently, when tasked with implementing functions that use non-standard naming (e.g., camelCase), an LLM trained on SGCR-processed data may rewrite function names to adhere to snake_case. This leads to mismatches during MBPP evaluation, where calling a function with its original non-standard name results in an “is not defined” error.
Figure 9 illustrates this issue: models trained on SGCR-processed data exhibit lower MBPP@1 and MBPP@10 scores compared to models trained on data processed only with syntax-error and linter-based filtering (Section 3.2.2). This performance drop stems from the naming mismatches described above, which obscure the model’s true code generation capabilities. Based on this finding, we concluded that MBPP is not a suitable benchmark for evaluating LLM code generation in our experiments, as its evaluation framework penalizes adherence to standard Python naming conventions. Therefore, we excluded MBPP from the benchmarks used in this study.
Figure 9: Comparison of MBPP@1 and MBPP@10 scores for models trained on linter-filtered data versus SGCR-rewritten data in a 50B-token continual pre-training ablation study. SGCR’s enforcement of snake_case naming conventions leads to lower scores due to mismatches with MBPP’s non-standard function names.
Appendix H Evaluations
In this section, we present the evaluation results of models trained through ablation experiments on code and math datasets, assessed across ten benchmarks encompassing code and mathematical downstream tasks.
H.1 Code ablation experiments results
As described in Section 3.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. In the following, we report the evaluation results for 13 code ablation experiments, with their relationships illustrated in Figure 10. Experiments exp1, exp8, exp9, and exp13 serve as baselines, utilizing only Python data extracted from existing open code corpora. The remaining experiments are conducted to construct the SwallowCode dataset. We evaluated performance using the following ten benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31].
Figure 10: Relationships between code ablation experiments.
Table 6: Performance across benchmarks in the-stack-v2-train-smol-ids Python subset ablation.
Experiment 1: the-stack-v2-train-smol-ids Python subset
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6659 0.5995 0.3354 0.9032 0.6294 0.4602 0.6019 0.3366 0.2994
20 0.3540 0.6567 0.6019 0.3360 0.9024 0.6238 0.4852 0.5898 0.3433 0.2890
30 0.3700 0.6588 0.6034 0.3377 0.9045 0.6263 0.5072 0.5939 0.3402 0.2951
40 0.3800 0.6618 0.6053 0.3380 0.9097 0.6341 0.5011 0.6016 0.3659 0.3079
50 0.3700 0.6679 0.6054 0.3350 0.9045 0.6340 0.5027 0.6091 0.3689 0.3140
Table 7: Performance across benchmarks in syntax-error-free data ablation from Experiment 1.
Experiment 2: Syntax-error-free data from Experiment 1
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3560 0.6675 0.6015 0.3385 0.9062 0.6321 0.4784 0.5881 0.3604 0.3030
20 0.3520 0.6635 0.6026 0.3364 0.9049 0.6252 0.4784 0.5781 0.3591 0.2970
30 0.3560 0.6637 0.6012 0.3375 0.9080 0.6313 0.5019 0.5950 0.3701 0.3122
40 0.3580 0.6679 0.6046 0.3346 0.9062 0.6330 0.5019 0.5998 0.3720 0.3067
50 0.3660 0.6694 0.6055 0.3340 0.9084 0.6325 0.5155 0.6044 0.3787 0.3110
Table 8: Performance across benchmarks in syntax-error and Pylint-filtered (score $\geq$ 7) data ablation from Experiment 2.
Experiment 3: Syntax-error and Pylint-filtered (score $\geq$ 7) data from Experiment 2
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3560 0.6628 0.6010 0.3340 0.9071 0.6235 0.4564 0.6007 0.3500 0.2811
20 0.3500 0.6613 0.6015 0.3361 0.9054 0.6237 0.4860 0.5838 0.3744 0.3207
30 0.3620 0.6596 0.6008 0.3359 0.9080 0.6307 0.4867 0.5921 0.3957 0.3134
40 0.3720 0.6650 0.6030 0.3352 0.9058 0.6326 0.4822 0.5990 0.3890 0.3341
50 0.3740 0.6677 0.6054 0.3291 0.9019 0.6327 0.4996 0.6145 0.3945 0.3317
Table 9: Performance across benchmarks in comment-language-restricted (English and Japanese) data ablation from Experiment 3.
Experiment 4: Comment-language-restricted (English and Japanese) data from Experiment 3
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6713 0.5988 0.3329 0.9054 0.6312 0.4708 0.5953 0.3549 0.3079
20 0.3520 0.6601 0.6011 0.3306 0.9067 0.6250 0.4898 0.5802 0.3689 0.3134
30 0.3680 0.6596 0.6047 0.3365 0.9118 0.6301 0.4989 0.5890 0.3768 0.3299
40 0.3660 0.6671 0.6049 0.3363 0.9071 0.6333 0.5155 0.6024 0.3756 0.3348
50 0.3700 0.6703 0.6061 0.3357 0.9101 0.6347 0.5133 0.6036 0.3841 0.3354
Table 10: Performance across benchmarks in SGCR-rewritten data ablation from Experiment 3.
Experiment 5: SGCR-rewritten data from Experiment 3
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3560 0.6689 0.5996 0.3295 0.9054 0.6256 0.4875 0.5991 0.4128 0.3652
20 0.3460 0.6610 0.6031 0.3352 0.9032 0.6262 0.4920 0.5801 0.4311 0.3805
30 0.3620 0.6637 0.6043 0.3378 0.9110 0.6269 0.5216 0.5984 0.4726 0.4293
40 0.3660 0.6645 0.6053 0.3372 0.9045 0.6328 0.4989 0.5945 0.4610 0.4128
50 0.3660 0.6667 0.6066 0.3325 0.9058 0.6352 0.5027 0.6065 0.4860 0.4110
Table 11: Performance across benchmarks in LLM-scored (score $\geq$ 6) data ablation from Experiment 3.
Experiment 6: LLM-scored (score $\geq$ 6) data from Experiment 3
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6679 0.6002 0.3277 0.9041 0.6280 0.4701 0.5976 0.3640 0.3134
20 0.3540 0.6593 0.6010 0.3358 0.9045 0.6249 0.4822 0.5810 0.3659 0.3110
30 0.3660 0.6594 0.6021 0.3398 0.9071 0.6226 0.5140 0.5893 0.3994 0.3396
40 0.3700 0.6636 0.6021 0.3370 0.9080 0.6300 0.5027 0.6019 0.4018 0.3329
50 0.3640 0.6684 0.6046 0.3353 0.9084 0.6324 0.5011 0.6090 0.3951 0.3402
Table 12: Performance across benchmarks in mixed (1:1) data ablation from Experiments 3 and 5.
Experiment 7: Mixed (1:1) data from Experiments 3 and 5
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3620 0.6660 0.5994 0.3293 0.9032 0.6242 0.4738 0.6156 0.3616 0.3061
20 0.3460 0.6585 0.6018 0.3297 0.9024 0.6293 0.4845 0.5809 0.3823 0.3427
30 0.3680 0.6611 0.6022 0.3384 0.9062 0.6241 0.5110 0.6045 0.3848 0.3427
40 0.3640 0.6666 0.6028 0.3327 0.9088 0.6323 0.5072 0.6056 0.4018 0.3634
50 0.3680 0.6695 0.6052 0.3320 0.9097 0.6300 0.5027 0.6051 0.4116 0.3573
Table 13: Performance across benchmarks in the-stack-v1 Python subset ablation.
Experiment 8: the-stack-v1 Python subset
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3660 0.6646 0.6033 0.3310 0.9028 0.6219 0.4784 0.5955 0.3244 0.2780
20 0.3500 0.6595 0.6018 0.3233 0.9037 0.6246 0.4701 0.5898 0.3220 0.2720
30 0.3640 0.6575 0.6014 0.3279 0.9071 0.6226 0.5057 0.5878 0.3244 0.2957
40 0.3680 0.6638 0.6029 0.3265 0.9067 0.6320 0.5004 0.5984 0.3445 0.3116
50 0.3620 0.6650 0.6053 0.3212 0.9084 0.6273 0.5080 0.5998 0.3561 0.3177
Table 14: Performance across benchmarks in codepartto-clean Python subset ablation.
Experiment 9: codepartto-clean Python subset
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3540 0.6651 0.6006 0.3221 0.9062 0.6295 0.4708 0.5875 0.3598 0.3128
20 0.3560 0.6556 0.6013 0.3358 0.9067 0.6289 0.4731 0.5870 0.3549 0.2927
30 0.3680 0.6570 0.6045 0.3390 0.9071 0.6290 0.4890 0.5976 0.3524 0.3201
40 0.3720 0.6613 0.6048 0.3352 0.9075 0.6300 0.4958 0.6108 0.3543 0.2988
50 0.3600 0.6638 0.6055 0.3321 0.9097 0.6273 0.5072 0.6139 0.3616 0.3134
Table 15: Performance across benchmarks in SGCR-rewritten data ablation from Experiment 1.
Experiment 10: SGCR-rewritten data from Experiment 1
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6667 0.5996 0.3325 0.9032 0.6164 0.4845 0.5959 0.4457 0.3518
20 0.3460 0.6592 0.6032 0.3324 0.9067 0.6231 0.4890 0.5655 0.4579 0.3994
30 0.3660 0.6585 0.6029 0.3379 0.9101 0.6176 0.5064 0.5855 0.4494 0.3884
40 0.3600 0.6650 0.6024 0.3339 0.9067 0.6284 0.5148 0.5967 0.4622 0.4067
50 0.3600 0.6687 0.6047 0.3337 0.9084 0.6317 0.5057 0.6041 0.4646 0.4067
Table 16: Performance across benchmarks in SCOR-rewritten data ablation from Experiment 5.
Experiment 11: SCOR-rewritten data from Experiment 5
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3580 0.6680 0.6006 0.3317 0.9067 0.6237 0.4655 0.6108 0.4567 0.4348
20 0.3500 0.6564 0.6026 0.3349 0.9084 0.6241 0.4981 0.5718 0.5384 0.4628
30 0.3620 0.6640 0.6023 0.3385 0.9054 0.6253 0.5095 0.5928 0.5256 0.4817
40 0.3640 0.6705 0.6041 0.3401 0.9088 0.6317 0.5095 0.5982 0.5226 0.4768
50 0.3700 0.6685 0.6055 0.3359 0.9114 0.6322 0.5110 0.6062 0.5396 0.4805
Table 17: Performance across benchmarks in mixed (90% Experiment 11, 10% Japanese-translated comments) data ablation.
Experiment 12: Mixed (90% Experiment 11, 10% Japanese-translated comments)
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6625 0.6020 0.3341 0.9054 0.6221 0.4738 0.5697 0.4799 0.4280
20 0.3500 0.6551 0.6021 0.3361 0.9058 0.6266 0.4943 0.5776 0.5165 0.4646
30 0.3640 0.6595 0.6034 0.3410 0.9080 0.6250 0.5011 0.6008 0.5110 0.4415
40 0.3640 0.6640 0.6022 0.3361 0.9054 0.6330 0.4898 0.6008 0.5299 0.4768
50 0.3600 0.6655 0.6057 0.3340 0.9080 0.6315 0.5072 0.6057 0.5329 0.4866
Table 18: Performance across benchmarks in Stack Edu Python subset ablation.
Experiment 13: Stack Edu Python subset
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU GSM8K BBH HumanEval HumanEval+
10 0.3640 0.6696 0.5986 0.3358 0.9037 0.6246 0.4761 0.6004 0.3470 0.2963
20 0.3520 0.6632 0.6021 0.3364 0.9067 0.6233 0.4898 0.5942 0.3537 0.2957
30 0.3660 0.6600 0.6024 0.3439 0.9097 0.6251 0.4989 0.5916 0.3713 0.2988
40 0.3700 0.6650 0.6033 0.3402 0.9067 0.6325 0.4958 0.6084 0.3701 0.3226
50 0.3740 0.6665 0.6061 0.3368 0.9062 0.6350 0.5087 0.6173 0.3695 0.3195
H.2 Math ablation experiments results
As described in Section 4.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. Below, we present the evaluation results for two math ablation experiments. Specifically, we adopted the following ten evaluation benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and MATH [27].
Table 19: Performance across benchmarks in finemath-4+ ablation.
Experiment 1: finemath-4+
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU HumanEval GSM8K BBH MATH
10 0.3700 0.6626 0.5990 0.3350 0.8985 0.6243 0.3439 0.4685 0.6057 0.1760
20 0.3720 0.6536 0.5963 0.3510 0.9032 0.6261 0.3622 0.5011 0.5896 0.2080
30 0.3700 0.6574 0.5999 0.3506 0.8998 0.6253 0.3561 0.5019 0.5971 0.2260
40 0.3720 0.6577 0.6024 0.3499 0.9049 0.6312 0.3701 0.5231 0.6054 0.2260
50 0.3740 0.6608 0.6001 0.3550 0.9058 0.6329 0.3561 0.5292 0.6166 0.2400
Table 20: Performance across benchmarks in finemath-4+ rewritten with Llama-3.3-70B-Instruct ablation.
Experiment 2: finemath-4+ rewriten(Llama-3.3-70B-Instruct)
Tokens (B) OpenBookQA TriviaQA HellaSwag SQuAD2.0 XWINO MMLU HumanEval GSM8K BBH MATH
10 0.3720 0.6643 0.5970 0.3443 0.9015 0.6343 0.3439 0.5603 0.5535 0.2480
20 0.3800 0.6580 0.5946 0.3428 0.8994 0.6293 0.3762 0.6156 0.5669 0.2860
30 0.3660 0.6618 0.5964 0.3470 0.9011 0.6298 0.3530 0.6262 0.6383 0.3040
40 0.3700 0.6610 0.5973 0.3535 0.9088 0.6358 0.3738 0.6422 0.6237 0.3100
50 0.3800 0.6637 0.5972 0.3537 0.9045 0.6337 0.3683 0.6535 0.6414 0.3160